OpenStudy (anonymous):

How to Solve (g+4)-3g=1+g?

OpenStudy (anonymous):

There really arent any "steps" g+4-3g = 1 + g subtract a g from both sides 4 - 3g = 1 subtract 4 from both sides -3g = -3 g = 1

OpenStudy (anonymous):

Oh okay thanks.

OpenStudy (anonymous):

yw!